OpenStudy (anonymous):

Can u guys check my answer? What is the approximate volume of a sphere with a radius of 12 ft? Use 3.14 to approximate pi and express your answer in hundredths. 77234.55998 ft?

OpenStudy (anonymous):

@MrNood

OpenStudy (phi):

what formula did you use ?

OpenStudy (anonymous):

4/3*3.14*r^3

OpenStudy (phi):

and r is 12 so figure out 4*3.14*12*12*12/3

OpenStudy (anonymous):

I already did this.

OpenStudy (phi):

do it again

OpenStudy (anonymous):

the formula is 4/3*3.14*r^3

OpenStudy (phi):

yes.

OpenStudy (anonymous):

i got 7264.56001

OpenStudy (phi):

you have the correct formula, but you did something wrong when doing the arithmetic if you use 3.14 for pi, you will only get 2 decimals in that answer.

OpenStudy (anonymous):

what did i do wrong?

OpenStudy (phi):

notice that 4*3.14*12*12*12/3 the last 12/3 is 4, so the problem is 4*3.14*12*12*4 which you can write as 4*4 * 12*12 * 3.14 or 16* 144 * 3.14

OpenStudy (phi):

First get the correct answer. then we can figure out what you did wrong.

OpenStudy (anonymous):

7234.56

OpenStudy (phi):

yes. that looks good. now try 4*3.14*12*12*12/3 on you calculator. you should get 7234.56

OpenStudy (anonymous):

but y do u have to divide by 3?

OpenStudy (anonymous):

and i did get 7234.56

OpenStudy (phi):

I wrote 4/3 * 12*12*12*3.14 as 4*12*12*12*3.14/3 if you write it as a fraction: \[ \frac{4}{3} \cdot \frac{3.14 \cdot 12^3 }{1} \] and multiply top times top and bottom times bottom: \[ \frac{3.14 \cdot 12^3 }{3} \] and that means you can first figure out the top, and then divide by 3

OpenStudy (phi):

If you play with numbers, you learn these ideas

OpenStudy (anonymous):

ok thx

OpenStudy (phi):

* I forgot the 4 up top \[ \frac{4 \cdot 3.14 \cdot 12^3 }{3}\]

OpenStudy (phi):

Can you figure out what you did to get 77234.55998 ?

OpenStudy (phi):

because you don't want to do that

OpenStudy (anonymous):

ohh ok i get i know thx a lot.